OpenStudy (anonymous):

answer this question and get a medal A grocer wants to make a 10-pound mixture of cashews and peanuts that he can sell for $3.29 per pound. If cashews cost $5.60 per pound and peanuts cost $2.30 per pound, how many pounds of each must he mix?

OpenStudy (radar):

Lets use algebra solve this

OpenStudy (anonymous):

k

OpenStudy (anonymous):

You need to set up a system of equations and then solve. See if you can solve this system of equations. x+y=10 2.30x+5.60y=10(3.29) Where x = pounds of peanuts and y = pounds of cashews

OpenStudy (radar):

Let x = pound of peanuts and then let y= pounds of Cashers wait, I see NinjaDevl is going to assist you, I stand by.

OpenStudy (anonymous):

alright ill try

OpenStudy (radar):

*NinjaDevo Not NinjaDevl

OpenStudy (anonymous):

Ok, let me know if you get stuck. Btw, sorry, didn't mean to steal the question from you @radar

OpenStudy (anonymous):

alright

OpenStudy (radar):

No problem, I was going to set it up exactly as you have, then use substitution to solve the system.

OpenStudy (radar):

You still there @tylerman866

OpenStudy (anonymous):

ya trying to work it out. sorry suck at math it doesn't come easily to me

OpenStudy (anonymous):

it = 3.29 right?

OpenStudy (radar):

It is a system of equations, 2 equations 2 unknowns. It can be solved. Have you got into "system of linear equations" yet?

OpenStudy (radar):

The 3.29 is what he intends to sell the mixture per pound.

OpenStudy (anonymous):

no i haven't gotten into it yet

OpenStudy (anonymous):

so what would i do now that i figured out he is selling it for 3.29 a pound

OpenStudy (anonymous):

I'll let you take this over from here radar. :)

OpenStudy (anonymous):

hey thanks for the help NinjaDevo

OpenStudy (radar):

NinjaDevo provided two equations: X + Y = 10 That is the amount of pounds he wants for the cashew/peanuts mix. 2.30X + 5.60Y = 10 * 3.29 is the money the cost of the peanuts and cashews (x pounds) and y pounds. Mission: Express 2nd equation using only one variable choose the variable (x or y) which you want to use.

OpenStudy (radar):

That way we will be working with one equation with one variable..........which one did you choose?

OpenStudy (anonymous):

x

OpenStudy (radar):

The choice is really arbitrary, you canwork with either one, ah x is the choice the peanuts.

OpenStudy (radar):

To work just with x we need to express y in terms of x . Using the first equation X + Y = 10 Solving for y would give us, y = 10 - x. Do you follow the algebra there?

OpenStudy (anonymous):

ya kinda

OpenStudy (anonymous):

sorta?

OpenStudy (radar):

We just subtracted y from both sides....Thats legal Mission continues. Taking the 2nd equation (the big one) we now substitute for y, for y we can now simply use (10-x) Here goes: 2.30X + 5.60(10-x) = (3.29) (10) Note that we now have one equation with one unknown, the kinds you have been solving all along in algebra. Do you see how we have substituted for y?

OpenStudy (anonymous):

−32.9−2.30X5.60+1

OpenStudy (anonymous):

ya i dont got it i guess but i tried but thanks for the help

OpenStudy (radar):

Mission continues. we now do the operations as indicated by our new equation which has only x 2.30x + 56.0 - 5.6x =32.9 combining -3.3x + 56.0 = 32.9 subtracting 56 from both sides (to get x by itself) -3.3x = 32.9 - 56. = -23.1 dividing both sides by -3.3 x = 7 y then must equal 3 7 pounds of peanuts 3 pounds of cashews.

OpenStudy (radar):

You really didn't try, I thought you were needing help.

OpenStudy (anonymous):

so its ten pounds all together, or do you leave them seperated like that??

OpenStudy (radar):

Focus and practice and you will. Yes the 10 pounds was given by the problem, they wanted you to figure how many pounds of peanuts and how many pounds of cashews you would mix together to be able to sell the mix at 3.29 per pound.

OpenStudy (anonymous):

thank you so much!! are you a math teacher??

OpenStudy (radar):

No, a retired electronic engineer from the FAA

OpenStudy (anonymous):

damn that sounds complicated

OpenStudy (anonymous):

i guess i just need to apply myself and i could do something similar

OpenStudy (radar):

Good luck with your studies and good luck with whatever career field you so choose. You can do what you set your mind to.

OpenStudy (anonymous):

thanks